# An Ising Similarity Regression Model

### v1

@@ 4 Simulation Study
Table 1: MSE for the oracle, regularized, and unregularized estimators of the regression coefficients ($\textrm{MSE}_{\alpha}$) and main effect parameters ($\textrm{MSE}_{\theta}$) in the simulation study involving the Ising similarity regression model.
$\textrm{MSE}_{\alpha}$ $\textrm{MSE}_{\theta}$
$p$ $n$ Oracle Regularized Unregularized Oracle Regularized Unregularized
10 50 0.046 0.068 0.352 0.350 0.402 0.896
100 0.022 0.044 0.147 0.167 0.207 0.370
200 0.011 0.028 0.068 0.082 0.111 0.172
400 0.005 0.016 0.033 0.038 0.057 0.080
25 50 0.006 0.015 0.026 0.298 0.391 0.546
100 0.003 0.007 0.012 0.137 0.184 0.242
200 0.001 0.003 0.006 0.068 0.088 0.118
400 0.001 0.001 0.003 0.033 0.041 0.057
50 50 0.002 0.006 0.008 0.550 0.788 0.737
100 0.001 0.002 0.004 0.120 0.156 0.205
200 $<$0.001 0.001 0.002 0.053 0.067 0.093
400 $<$0.001 $<$0.001 0.001 0.027 0.033 0.047
100 50 0.001 0.005 0.003 2.314 1.551 2.444
100 $<$0.001 0.001 0.001 0.500 0.559 0.567
200 $<$0.001 $<$0.001 0.001 0.099 0.111 0.132
400 $<$0.001 $<$0.001 $<$0.001 0.033 0.034 0.049
200 50 $<$0.001 0.005 0.001 1.971 2.180 2.073
100 $<$0.001 $<$0.001 0.001 0.290 0.350 0.343
200 $<$0.001 $<$0.001 $<$0.001 0.073 0.076 0.102
400 $<$0.001 $<$0.001 $<$0.001 0.030 0.031 0.044
@@
Table 1 shows that the MSEs for all three estimators of the regression coefficients $\bm{\alpha}$ decreased as the sample size $n$ and the number of responses $p$ increased. The MSEs of the main effect parameters exhibited a clear decreasing trend when $n$ increased, but no obvious patterns when $p$ increased; this is to be expected given the main effect parameter $\theta_{jj}$ was specific to each of the responses in the Ising similarity regression model for $j=1,\cdots,p$. Among the three estimators, the MSEs for the regularized estimators of the regression coefficients and main effect parameters were typically much smaller than the unregularized estimators, and in fact similar to the oracle estimators especially when $n$ and $p$ were large. This is consistent with the notion that the regularized estimators take advantage of the underlying sparsity in the model, resulting in better overall estimation performance than the unregularized estimators.
@@
Table 2: TPR and FPR for the regularized estimators of the regression coefficients in the simulation study involving the Ising similarity regression model.
True Positive Rate (TPR) False Positive Rate (FPR)
@@
Turning to the model selection performance of the regularized pseudo-likelihood estimator (Table 2), while there was slight underfitting as reflected by the comparably low TPR for the case of smallest $p=10$, both the TPR and FPR tended to one and zero, respectively, as $n$ and $p$ increased. This is consistent with Theorem 2(b), empirically demonstrating that the proposed method was able to recover the underlying sparsity in the Ising similarity regression model, while also providing empirical evidence for the use of the cross-validation approach in selecting $\lambda$.
@@
We applied the Ising similarity regression model (2.4) to roll call voting data from the U.S. Senate as part of the 117-th Congress, covering the period from 6 January 2021 to 20 May 2021 (date of data collection). Roll call voting data has previously been studied using a variety of statistical techniques including via undirected graphical models (Banerjee, El Ghaoui, and d’ Aspremont, 2008). Here, we use our model to study how voting associations between senators are associated with their similarities in various demographic attributes and social network profiles. The dataset was obtained from the U.S. Senate’s website (https://www.senate.gov/). It consists of the voting records on 199 bills by 100 senators of the 117-th Congress. The votes were coded as one for ‘Yea’ and zero for ‘Nay’, thus producing multivariate binary vectors $\bm{y}_{i}$ with elements $y_{ij}$ representing the binary vote on the $i$-th bill by the $j$-th senator. Similar to Banerjee, El Ghaoui, and d’ Aspremont (2008) and Guo et al. (2010), all bills were treated as independent. Furthermore, following Guo et al. (2010), we removed procedural votes with the ‘Yea/Nay’ proportion falling outside of the interval [0.3,0.7], as these were uncontroversial bills that did not reveal political dynamics in the U.S. Senate. The dataset also contained a small number of missing values since not all senators voted on every bill. Specifically, 1.44% of all votes were missing, and we chose to impute these with the majority vote of the senator’s party on the same bill. Note even though we used party to construct one of the similarity matrices later on in our model, we do not believe this form of imputation impacted our analysis to any great degree, given the percentage of the missing data was very small. The final dataset analyzed thus consisted of $n=138$ bills voted by $p=100$ senators.
@@
We obtained several attributes for each senator, including their state, political party, class, and age from the U.S. Senate website, along with gender and occupation from Wikipedia. Additionally, each senator’s Twitter handle was obtained from the ‘us-senate’ GitHub project of ‘CivilServiceUSA’, and used to compute the number of tweets and number of followers for each senator. Each of these attributes was then converted into a similarity matrix $\bm{W}_{k}$, where for qualitative variables (state, party, class and gender) we set $w_{jj^{\prime}}^{(k)}=1$ if the $j$-th and $j^{\prime}$-th senators are in the same category and zero otherwise, while for quantitative variables (age, number of tweets, number of Twitter followers) we used $w_{jj^{\prime}}^{(k)}=\exp(-|z_{jk}-z_{j^{\prime}k}|^{2})$. For the occupation, given there was such a large number of categories, we considered the seven most common occupations in the sample (lawyer, executive, businessman, farmer, army, teacher, professor) and constructed a separate similarity matrix for each of these occupations. That is, $w_{jj^{\prime}}^{(k)}=1$ if the $j$-th and $j^{\prime}$-th senators have the same specific occupation and zero otherwise. Finally, we constructed a symmetric adjacency matrix summarizing the Twitter follower-followee relationship among the 100 senators, such that $w_{jj^{\prime}}^{(k)}=1$ if the $j$-th senator follows the $j^{\prime}$-th senator on Twitter or vice versa, and zero otherwise. Combined, the above resulted in a total of $K=15$ similarity matrices.
@@
We fitted the proposed Ising similarity regression model using the regularized pseudo-likelihood estimator based on the adaptive lasso penalty to identify and quantify truly important attributes driving the voting associations between senators. As in the simulations of Section 4, the tuning parameter $\lambda$ was chosen using ten-fold cross-validation where the groupings were done at the bill level. After performing variable selection and obtained the estimated non-zero regression coefficients, we constructed 95% Wald confidence intervals for each coefficient based on standard errors obtained from the empirical sandwich covariance matrix; see Section LABEL:sec:additional_results_rollcall of the supplementary material for details of the empirical sandwich covariance matrix calculation.
@@
Table 3 reports the estimated regression coefficients for the $K=15$ similarity matrices along with their corresponding 95% Wald confidence intervals. The estimation results for the main effect parameters can be found in the supplementary material. Using the adaptive lasso penalty, only 7 of the 15 regression coefficients were estimated to be non-zero. Of these, unsurprisingly there was statistically clear evidence that senators from the same state and/or party were more likely to vote similarly on the bills (analogous state and party effects were found in Banerjee, El Ghaoui, and d’ Aspremont, 2008; Guo et al., 2010, among others), although the former exhibited a much stronger effect. Most occupations were found to be uninformative for the conditional dependence relationships between senator voting patterns, except for businessman and lawyer, although it should be noted that the confidence interval corresponding to the effect of lawyer contained zero. The presence of a positive effect for the businessman similarity matrix could be attributed to senators who were businessmen tending to vote similarly on bills related to the economy.
@@
There was statistically clear evidence of a positive association between the Twitter follower-followee relationship adjacency matrix and senator voting patterns. A possible explanation for this might be that one senator who followed the other senator on Twitter had more exposure to their advocated ideologies, and hence was more likely to vote similarly. It is also possible that the senator who followed the other senator on Twitter already agreed with their ideologies in the first place, while their interactions on Twitter further reinforced such agreement, leading to positive associations in their voting patterns. The similarity in terms of senators’ popularity on Twitter, as measured by the number of followers, did not have any effect on the association between senators’ votes. Interestingly, although the effect of the similarity matrix associated with senators’ number of tweets was found to be negative, when we ran a separate analysis based on fitting an Ising similarity regression model with only this similarity matrix, its associated coefficient became positive: $\hat{\alpha}_{\textrm{Tweets}}=0.083$ with 95% confidence interval being $(0.080,0.087)$. Therefore, we conjecture the negative coefficient found in Table 3 was likely due to a large amount of information contained within this similarity matrix that could be explained by the other similarity matrices i.e., a form of matrix multicollinearity.
@@
Figure 1 presents graphs of the weighted similarity/adjacency matrices $\hat{\alpha}_{k}\bm{W}_{k}$ for selected attributes, together with the estimated interaction matrix $\hat{\bm{\Theta}}$ describing the voting associations for a subset of 20 senators, where $\hat{\bm{\Theta}}=\sum_{j=1}^{100}\hat{\theta}_{jj}\bm{\Delta}_{jj}+\sum_{k=1}^{15}\hat{\alpha}_{k}\bm{W}_{k}$ based on equation (2.3), and $\hat{\alpha}_{k}$ and $\hat{\theta}_{jj}$ are the regularized pseudo-likelihood estimators of the regression coefficients and the main effect parameters, respectively. The choice of senators to be included was made by beginning with an empty set, and sequentially adding pairs of senators who had the largest off-diagonal elements $\hat{\theta}_{jj^{\prime}}$ in $\hat{\bm{\Theta}}$ until the set contained the top 20 senators. To further improve clarity of the graph based on the estimated $\hat{\bm{\Theta}}$, we removed all edges with $\hat{\theta}_{jj^{\prime}}\leq 0.1297$ (the median of $\{\hat{\theta}_{jj^{\prime}}:j\neq j^{\prime}\}$); this explains the visually denser graph of Twitter follower-followee relationship compared to that of $\hat{\bm{\Theta}}$ in Figure 1. It can be seen that the estimated strong voting associations (i.e., large estimated values of $\theta_{jj^{\prime}}$) between senators were being driven heavily not only by the similarities in senators’ state and party affiliations, but also their relationships on Twitter as indicated by the red-colored edges in the graph of $\hat{\bm{\Theta}}$. Indeed, the fitted Ising similarity regression model yielded a log pseudo-likelihood of -1125.28, compared to -9441.01 for a null model with only the main effect parameters. This corresponds to a pseudo $R^{2}$ (Cohen et al., 2013) of approximately 0.88, indicating the selected model provided a much better fit than the null model through the inclusion of the available similarity/adjacency matrices $\bm{W}_{k}$ to explain the dependence structure.
@@ 6 Conclusion
In this article, we developed an Ising similarity regression model to study the effect of predictor information, arising in the form of similarity matrices, on conditional dependence relationships among binary responses. A computationally efficient regularized pseudo-likelihood estimator using the adaptive lasso penalty was proposed, which we demonstrated to be estimation and model selection consistent under a general setting where the number of similarity matrices $K$ and responses $p$ grow with sample size $n$. Simulations demonstrated the strong finite sample point estimation and selection performance of the proposed estimator. Applying the Ising similarity regression model to the U.S. Senate roll call voting data not only identified well-documented state and party effects in driving senator voting patterns, but also established new insights into the importance of senators’ similarities in businessman occupation and social network relationships on their voting dependence. These findings are new and differ from those obtained from the graphical structure analyses using standard Ising models (Guo et al., 2010, 2015). In particular, while graphical structure analyses aim to reveal the dependency structures among responses (senators), our method seeks to recover and quantify how different auxiliary information (senator’s attributes) influence this structure. While various methods have been proposed to address the former problem, the Ising similarity regression model is one of the first to have been developed for addressing the latter.
@@
A logical next step would be to extend the proposed model to incorporate predictor information from both observations and responses, thus forming a sort of “double Ising similarity regression model”. In particular, this may involve accounting for spatial and/or temporal dependence between observations, where similarity matrices could then be built up based on some knowledge of spatial or temporal distance (e.g., Bonat and Jørgensen, 2016). Furthermore, establishing asymptotic normality results for the regularized pseudo-likelihood estimator would serve as a useful basis for inference, and provide some theoretical justification for our proposed empirical sandwich-based standard errors. Also, other penalty functions could be used to replace the adaptive lasso penalty for similarity selection, including the use of group or fusion penalties when the set of similarity matrices exhibits some sort of hierarchy or ordering (Hui, Müller, and Welsh, 2017a; Zhang et al., 2023). It would be useful to consider similarity matrices $\bm{W}_{k}^{(i)}$ that are heterogeneous across $i=1,\cdots,n$; this would correspond to extending the Ising similarity regression model to allow for heterogeneous interaction matrices $\bm{\Theta}^{(i)}$ for different binary response vectors $\bm{y}_{i}$ (see for instance Zou et al., 2022, in the context of covariance matrices of continuous response vectors). Finally, while this work has mainly focused on multivariate binary responses using the Ising model, similar modeling idea could be extended to quadratic exponential families (Gourieroux, Monfort, and Trognon, 1984) to allow for other response types.
@@
The Supplementary Material contains sample versions of Conditions 1 and 3 and proofs of the theorems, as well as additional details of application to the U.S. Senate roll call voting data.

### v2

@@ 4 Simulation Study
Table 1: MSE for the Oracle, Regularized, Lasso, and Unregularized estimators of the regression coefficients ($\textrm{MSE}_{\alpha}$) and main effect parameters ($\textrm{MSE}_{\theta}$) in the simulation study involving the Ising similarity regression model. $\textrm{MSE}_{\alpha}$ is multiplied by 1000 for clarity.
$1000\times\textrm{MSE}_{\alpha}$ $\textrm{MSE}_{\theta}$
$p$ $n$ Oracle Regularized Lasso Unregularized Oracle Regularized Lasso Unregularized
10 50 45.791 68.420 37.524 351.622 0.350 0.402 0.346 0.896
100 21.544 44.121 28.569 147.225 0.167 0.207 0.193 0.370
200 10.755 27.506 19.048 68.262 0.082 0.111 0.099 0.172
400 5.390 16.298 11.717 33.021 0.038 0.057 0.052 0.080
25 50 5.858 14.717 13.474 25.985 0.298 0.391 0.351 0.546
100 2.846 6.817 6.992 12.089 0.137 0.184 0.171 0.242
200 1.434 3.025 3.546 5.996 0.068 0.088 0.087 0.118
400 0.710 1.356 1.780 2.950 0.033 0.041 0.043 0.057
50 50 2.100 5.901 5.908 8.368 0.550 0.788 0.719 0.737
100 0.969 1.924 2.537 3.967 0.120 0.156 0.166 0.205
200 0.469 0.860 1.266 1.935 0.053 0.067 0.074 0.093
400 0.232 0.386 0.625 0.939 0.027 0.033 0.038 0.047
100 50 0.765 4.517 3.282 3.065 2.314 1.551 2.201 2.444
100 0.362 1.390 1.609 1.441 0.500 0.559 0.512 0.567
200 0.175 0.367 0.510 0.709 0.099 0.111 0.113 0.132
400 0.089 0.119 0.227 0.355 0.033 0.034 0.041 0.049
200 50 0.351 5.493 2.406 1.282 1.971 2.180 2.420 2.073
100 0.169 0.376 0.422 0.614 0.290 0.350 0.395 0.343
200 0.079 0.108 0.197 0.295 0.073 0.076 0.089 0.102
400 0.039 0.061 0.096 0.148 0.030 0.031 0.035 0.044
@@
Table 1 shows the $\textrm{MSE}_{\alpha}$ for all four estimators decrease as $n$ and $p$ increase, while their $\textrm{MSE}_{\theta}$ exhibit a clear decreasing trend when $n$ increases but no obvious patterns when $p$ increases; this is to be expected given the main effect parameter $\theta_{jj}$ is specific to each of the responses in the Ising similarity regression model for $j=1,\cdots,p$. The $\textrm{MSE}_{\alpha}$ and $\textrm{MSE}_{\theta}$ for the proposed regularized pseudo-likelihood estimators are typically much smaller than the unregularized estimators, and in fact similar to the oracle estimators especially when $n$ and $p$ are large. This is consistent with the notion that the proposed estimators take advantage of the underlying sparsity in the model, resulting in better overall estimation performance than the unregularized estimators. Furthermore, the MSEs of the proposed estimators tend to be smaller than the lasso-regularized estimators when $n$ and $p$ are large.
@@
Table 2: TPR and FPR for the Regularized (top panel) and Lasso (bottom panel) estimators of the regression coefficients in the simulation study involving the Ising similarity regression model.
Regularized
TPR FPR
@@
Lasso
TPR FPR
$p\backslash n$ 50 100 200 400 50 100 200 400
10 0.307 0.500 0.711 0.888 0.124 0.185 0.237 0.290
25 0.859 0.980 1 1 0.298 0.356 0.392 0.395
50 0.982 0.999 1 1 0.403 0.412 0.426 0.428
100 0.999 1 1 1 0.294 0.327 0.405 0.429
200 0.996 1 1 1 0.254 0.354 0.392 0.433
Turning to the model selection performance of the proposed estimator in Table 2, while there is slight underfitting as reflected by the comparably low TPR for the case of smallest $p=10$, both the TPR and FPR tend to one and zero, respectively, as $n$ and $p$ increase. This is consistent with Theorem 2(b), empirically demonstrating that the proposed method is able to recover the underlying sparsity in the Ising similarity regression model. The lasso-regularized estimator suffers from overfitting as seen from its FPR not converging towards zero, thus supporting the use of the adaptive lasso penalty in our proposed estimator.
Results from comparing our ten-fold cross-validation approach to the use of AIC and BIC criteria for choosing the tuning parameter are provided in Table LABEL:table:simulation_tpr_fpr_aic_bic of the supplementary material. Overall, results show the BIC performs similarly well as our cross-validation approach, while the AIC suffers from clear overfitting when $p$ is small due to its weaker model complexity penalty. These results provide empirical support for using the cross-validation approach to choose $\lambda$.
Next, we compare the proposed estimator to other traditional estimators of the Ising model in the literature that involve direct estimation of the Ising model interaction matrix $\bm{\Theta}$, without considering the similarity matrices $\bm{W}_{k}$ (e.g., Section 3 of Höfling and Tibshirani, 2009). From Table LABEL:table:simulation_fsnorm in the supplementary material, we see that the proposed estimator greatly outperforms the traditional Ising model estimators in recovering the true $\bm{\Theta}$ matrix, since it incorporates the additional information from the similarity measures.
Finally, we conduct simulation studies with varying number of similarity matrices $K\in\{10,20,40,80,200\}$ while keeping the same number of truly non-zero regression coefficients $K_{0}=5$, as a further investigation on the effect of $K$ on the empirical performance of various estimators. Unsurprisingly, the estimation and model selection performance of all methods decline as more irrelevant similarity matrices are being added, noting the proposed estimator still performs reasonably well relative to other estimators under different settings of $K$; see Tables LABEL:table:simulation_mse_varying_KLABEL:table:simulation_fsnorm_varying_K in the supplementary material for full results.
@@
We apply the Ising similarity regression model (2.4) to roll call voting data from the U.S. Senate as part of the 117-th Congress, covering the period from 6 January 2021 to 20 May 2021 (date of data collection). Roll call voting data has previously been studied using a variety of statistical techniques including undirected graphical models (Banerjee, El Ghaoui, and d’ Aspremont, 2008). Here, we use our model to study how voting associations between senators are associated with their similarities in various demographic attributes and social network profiles. The dataset is obtained from the U.S. Senate’s website (https://www.senate.gov/), which originally consists of binary voting records, coded as one for ‘Yea’ and zero for ‘Nay’, on 199 bills by 100 senators of the 117-th Congress. After performing some preliminary data wrangling procedures (see Section LABEL:sec:additional_results_rollcall of the supplementary material), the final dataset analyzed consists of $n=138$ bills voted by $p=100$ senators, where all bills are treated as independent (e.g., Banerjee, El Ghaoui, and d’ Aspremont, 2008; Guo et al., 2010).
@@
We obtain several attributes for each senator, including their state, political party, class, and age from the U.S. Senate website, along with gender and occupation from Wikipedia. Additionally, each senator’s Twitter handle is obtained from the ‘us-senate’ GitHub project of ‘CivilServiceUSA’, and used to compute the number of tweets and number of followers for each senator. Each of these attributes is then converted into a similarity matrix $\bm{W}_{k}$ following the procedure described in Section 2.1, depending on whether it is a qualitative (state, party, class, gender and occupation) or quantitative (age, number of tweets, number of Twitter followers) attribute. We refer the reader to Section LABEL:sec:additional_results_rollcall of the supplementary material for detailed construction of these similarity matrices, as well as a descriptive analysis for the above attributes and the binary votes of the senators. Note that a symmetric adjacency matrix $\bm{W}_{k}$ is also constructed to summarize the Twitter follower-followee relationship among the senators, such that $w_{jj^{\prime}}^{(k)}=1$ if the $j$-th senator follows the $j^{\prime}$-th senator on Twitter or vice versa, and zero otherwise. As a result, the analysis consists of $K=15$ similarity matrices.
@@
We fit the Ising similarity regression model using the regularized pseudo-likelihood estimator to identify and quantify truly important attributes driving the voting associations between senators, where $\lambda$ is selected using ten-fold cross-validation with the groupings being done at the bill level. After performing variable selection and obtaining the estimated non-zero regression coefficients, we construct 95% Wald confidence intervals for each coefficient based on standard errors from the empirical sandwich covariance matrix obtained by deriving the score and Hessian matrix of the log pseudo-likelihood for each observation with respect to the set of chosen similarity measures; see Section LABEL:sec:inference of the supplementary material for details of its derivation.
@@
Table 3 shows only 7 of the 15 regression coefficients are estimated to be non-zero using the adaptive lasso penalty; estimation results for the main effect parameters are given in the supplementary material. Of these, unsurprisingly there is statistically clear evidence that senators from the same state and/or party are more likely to vote similarly on the bills (analogous state and party effects are found in Banerjee, El Ghaoui, and d’ Aspremont, 2008; Guo et al., 2010, among others), although the former exhibits a much stronger effect. Most occupations are found to be uninformative for the conditional dependence relationships between senator voting patterns, except for businessman and lawyer, although the confidence interval corresponding to the effect of lawyer contains zero. The presence of a positive effect for the businessman similarity matrix could be attributed to senators who are businessmen tending to vote similarly on bills related to the economy.
@@
There is statistically clear evidence of a positive association between the Twitter follower-followee relationship adjacency matrix and senator voting patterns. A possible explanation is that one senator who follows the other senator on Twitter has more exposure to their advocated ideologies, and hence is more likely to vote similarly. It is also possible that the senator who follows the other senator on Twitter already agrees with their ideologies in the first place, while their interactions on Twitter further reinforce such agreement, leading to positive associations in their voting patterns. The similarity in terms of senators’ popularity on Twitter, as measured by the number of followers, does not have any effect on the association between senators’ votes. Interestingly, although the effect of the senators’ number of tweets is found to be negative, when we run a separate analysis based on fitting an Ising similarity regression model with only this similarity matrix, its associated coefficient becomes positive: $\hat{\alpha}_{\textrm{Tweets}}=0.083$ with 95% confidence interval being $(0.080,0.087)$. Therefore, the negative coefficient found in Table 3 is conjectured to be due to a large amount of information contained within this similarity matrix that could be explained by other similarity matrices i.e., a form of matrix multicollinearity.
@@
Figure 1 presents graphs of the weighted similarity/adjacency matrices $\hat{\alpha}_{k}\bm{W}_{k}$ for selected attributes, together with the estimated interaction matrix $\hat{\bm{\Theta}}$ describing the voting associations for a subset of 20 senators, where $\hat{\bm{\Theta}}=\sum_{j=1}^{100}\hat{\theta}_{jj}\bm{\Delta}_{jj}+\sum_{k=1}^{15}\hat{\alpha}_{k}\bm{W}_{k}$ based on equation (2.3), and $\hat{\alpha}_{k}$ and $\hat{\theta}_{jj}$ are the regularized pseudo-likelihood estimators. The choice of senators to be included is made by beginning with an empty set, and sequentially adding pairs of senators who have the largest off-diagonal elements $\hat{\theta}_{jj^{\prime}}$ in $\hat{\bm{\Theta}}$ until the set contains the top 20 senators. To further improve clarity of the graph based on the estimated $\hat{\bm{\Theta}}$, we remove all edges with $\hat{\theta}_{jj^{\prime}}\leq 0.1297$ (the median of $\{\hat{\theta}_{jj^{\prime}}:j\neq j^{\prime}\}$); this explains the visually denser graph of Twitter follower-followee relationship compared to that of $\hat{\bm{\Theta}}$ in Figure 1. We emphasize that the removal of edges here is purely to improve the visualization and interpretability of the graph, noting our focus is on similarity selection and not edge selection as discussed earlier in Sections 1 and 2.2. The corresponding graph without such removal of edges can be found in Figure LABEL:fig:dense_theta in the supplementary material, and provides qualitatively similar conclusions as above. It can be seen that the estimated strong voting associations (i.e., large estimated values of $\theta_{jj^{\prime}}$) between senators are being driven heavily not only by similarities in senators’ state and party affiliations, but also their relationships on Twitter as indicated by the red-colored edges in the graph of $\hat{\bm{\Theta}}$. Indeed, the fitted Ising similarity regression model yields a log pseudo-likelihood of -1125.28, compared to -9441.01 for a null model with only the main effect parameters. This corresponds to a pseudo $R^{2}$ (Cohen et al., 2013) of approximately 0.88, indicating the selected model provides a much better fit than the null model through the inclusion of the available $\bm{W}_{k}$ matrices to explain the dependence structure.
@@ 6 Conclusion
We develop an Ising similarity regression model to study the effect of predictor information, arising in the form of similarity matrices, on conditional dependence relationships among binary responses. A computationally efficient regularized pseudo-likelihood estimator using the adaptive lasso penalty is proposed, which we demonstrate to be estimation and model selection consistent under a general setting where the number of similarity matrices $K$ and responses $p$ grow with sample size $n$. Simulations demonstrate the strong finite sample point estimation and selection performance of the proposed estimator, especially compared with several traditional Ising model estimators in recovering the interaction matrix, as it incorporates additional information from relevant similarity matrices. The cross-validation approach for choosing the tuning parameter is shown to have comparable performance to the BIC in terms of model selection. Applying the proposed model to the U.S. Senate roll call voting data not only identifies well-documented state and party effects in driving senator voting patterns, but also establishes new insights into the importance of senators’ similarities in businessman occupation and social network relationships on their voting dependence. These findings are new and differ from those obtained from the graphical structure analyses using standard Ising models (Guo et al., 2010, 2015). In particular, while graphical structure analyses aim to reveal dependency structures among responses (senators) through edge selection, our method seeks to recover and quantify how different auxiliary information (senator’s attributes) influence this structure via similarity selection. While various methods have been developed for the former problem, our approach is one of the first to have been developed for addressing the latter.
@@
A logical next step would be to extend the proposed model to incorporate predictor information from both observations and responses, thus forming a sort of “double Ising similarity regression model”. This may involve relaxing the assumption of $\bm{y}_{1},\cdots,\bm{y}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}f(\cdot;\bm{\vartheta})$ to consider correlated data such as spatial and/or temporal data, where similarity matrices could be built up based on knowledge of spatial or temporal distance (e.g., Bonat and Jørgensen, 2016) to account for spatial or temporal dependence between observations. The assumption could also be relaxed by extending the Ising similarity regression model to allow for heterogeneous interaction matrices $\bm{\Theta}^{(i)}$ for different $\bm{y}_{i}$ vectors, e.g. by considering $\bm{W}_{k}^{(i)}$ that are heterogeneous across $i=1,\cdots,n$; see for instance, Zou et al. (2022) who model heterogeneous covariance matrices of continuous response vectors based on linear combination of heterogeneous similarity matrices.
While this article focuses on the estimation and selection of the regression coefficients $\bm{\alpha}$, future research would involve rigorous theoretical study of the inferential component of our method e.g., by establishing asymptotic normality results for the regularized pseudo-likelihood estimator. Given the comparable model selection performance of the BIC to our cross-validation method, it would also be useful to extend this investigation to consider other types of information criterion in selecting $\lambda$ (e.g., Zhang, Li, and Tsai, 2010; Fan and Tang, 2013). On a related note, other penalty functions could be used to replace the adaptive lasso penalty for similarity selection, including the use of group or fusion penalties when the set of similarity matrices exhibits some sort of hierarchy or ordering (Hui, Müller, and Welsh, 2017a; Zhang et al., 2023). Finally, while this work focuses on binary responses using the Ising model, similar modeling idea could be extended to quadratic exponential families (Gourieroux, Monfort, and Trognon, 1984) to allow for other response types.
@@
The Supplementary Material contains sample versions of Conditions 1 and 3, proofs of the theorems, inference method, additional simulation results, along with supplementary details of application to the U.S. Senate roll call voting data, as well as an additional application to the Scotland Carabidae ground beetle data.
Acknowledgements
Zhi Yang Tho was supported by an Australian Government Research Training Program scholarship. Francis KC Hui was supported by an Australian Research Council Discovery Project DP230101908. Thanks to Alan Welsh for useful discussions.